# Opti-Acoustic Semantic SLAM with Unknown Objects in Underwater Environments

### v1

@@ I Introduction
Figure 2: The pipeline receives sensor inputs from a multibeam sonar, a monocular camera, an IMU, a DVL, and a pressure sensor. Given those sensor inputs, the system does the following: (1) Segment objects from optical images and extract features for each segmentation mask. The dimension of these extracted features is dependent on the corresponding object’s size in the image. These features are projected into a fixed dimension for the data association process. (2) Given the segmentation mask, its pixel centroid is used to find a corresponding range from sonar returns and estimate a 3D position of the object. (3) In parallel, sensor readings from IMU, DVL, and pressure sensor are used to estimate odometry. Outputs from (1)-(3) are used to build a factor graph, and we optimize it via iSAM2 to estimate maximum a posteriori (MAP). Lastly, we obtain map and trajectory estimates as an output.
@@ II-B Opti-Acoustic Correspondences
Unfortunately, each sensor modality has a degeneracy stemming from the loss of the elevation angle for sonar, and the loss of the range measurement for monocular cameras. Some works seek to address the degeneracy directly for sonar-only mapping by learning the elevation angle [21], through the fusion of orthogonally mounted sonars [22], or through carefully handling the degeneracy in the back-end [23]. Stereo or multi-view for optical-camera only mapping has been explored as well, but suffers from issues with lighting conditions, reflections, caustics, and featureless areas such as plain walls [7]. However, combining the two sensor modalities to address the degeneracies of each sensor modality is likely the best way to fully utilize their complimentary geometries.
@@
Jang et al. [28] attempt to learn opti-acoustic feature correspondences through the use of a convolutional neural network, while Jang et al. [29] try to calculate opti-acoustic feature matching by applying a style-transfer method and through integration with SuperGlue for multi-session or multi-vehicle mapping. Lanthier et al. [7] combine a stereo camera with sonar and infrared sensors to perform mapping, but note that stereo cameras struggle with obtaining range measurements for underwater vision due to uniform and featureless regions, and difficult lighting effects such as reflections and caustics. Yang et al. [30] use sonar to provide an absolute scale to monocular odometry. Kim et al. [31] separately perform visual SLAM and space carving sonar in two different mapping modules, before combining the two for dense reconstruction. Rahman et al. [32] combine sonar and monocular camera data in a SLAM framework they refer to as SVIN2, but their mapping technique uses point features rather than objects.
The only work that performs fusion at the object level as required by our method, rather than at the feature level, is by Raaj et al. [33], who use a monocular camera and sonar in a particle filter to track a single object.
@@ III-B Opti-acoustic Fusion for 3D Projection
Refer to captionRefer to captionRefer to captionRefer to captionptRaw ImageptSegmented ObjectptRaw SonarptEnhanced for VisualizationptTrue RangeptErroneous Range
@@ III-C Semantic SLAM
For each existing landmark that is determined to meet the similarity threshold, the Mahalanobis distance is calculated between the observed landmark and the existing landmark following the method described by Kaess et al. [47]. The distance follows a six degree-of-freedom chi-square distribution, and thus we use the test below where $l$ is the new landmark, $l_{j}$ is the existing landmark, and $\Sigma$ is the recovered covariance matrix.
@@
Figure 5: (Top): The outdoor tank used for data collection exhibits challenging lighting conditions, including reflections (left) and caustics (right). (Bottom): Selected frames from the vehicle’s monocular camera. These images demonstrate the caustics (ripple lighting effect due to surface waves), reflections, and sudden lighting changes make processing underwater imagery particularly difficult compared to terrestrial scenes. Feature-based methods fail to maintain tracking due to the dynamic textures; our method uses objects instead and proves to be robust to such lighting effects.
@@ IV Datasets
Figure 6: Comparison results of trajectory estimates, where the ’O’ and ’X’ markers denote the start and end points of each lap, respectively. The start and end points of the true trajectory are the same location, but drift causes the vehicle’s odometric estimate to deviate over time. (a) The two lap trajectory (Run 3) estimated by our patch averaging method (PA) can reduce drift errors from the odometry trajectory. (b) A similar comparison over four laps (Run 4), highlighting corrections made by our method against the odometry.
Refer to captionRefer to captionptLobster cageptSeaweedptHand netptTowfishptTireptPipe
@@
IV-A Outdoor Tank
Data was collected in an outdoor tank on a sunny day with a modified BlueROV2 that has an acoustic array for the iLBL-USBL measurements attached to the top of the vehicle. The vehicle also has a Bar30 pressure sensor for depth measurements, a Sony Exmor IMX322 low-light camera with an 80-degree horizontal field-of-view and a 64-degree vertical field-of-view, and an Oculus m750d multibeam sonar operating at 750kHz frequency with a 5 Hz update rate. In this mode, the sonar has a range resolution of 4 mm, an angular resolution of 1 degree, a horizontal field-of-view of 130 degrees, and a vertical aperture of 20 degrees. The relatively large vertical aperture, which cannot be resolved in the received data, means that the use of the camera for estimating the elevation angle of objects is crucial.
@@ IV-A Outdoor Tank
Since magnetometer-based heading measurements are adversely affected by the magnetic interference of the tank walls and other nearby equipment, to get a more accurate ground truth trajectory, the vehicle heading is directly inferred from beamforming on the measurements from the vehicle’s acoustic array by solving for vehicle heading $\theta_{k}$
@@
TABLE I: Summary of Indoor Datasets
@@
1
$1$ small hand net, $1$ large hand net, $2$ sets of stairs, $2$ recycling bins, $2$ types of towfish, $2$ trash bins, $2$ paddles
$2$
2
$1$ model submarine, $1$ green seaweed patch, $1$ pipe, $1$ towfish, $1$ star-shaped lobster cage, $1$ tire, $1$ small hand net
$2$
3
$2$ round lobster cages, $1$ green seaweed patch, $1$ blue seaweed patch, $2$ tires, $1$ small hand net, $2$ types of towfish, $2$ pipes with anchors, $1$ star-shaped lobster cage
$2$
4
$2$ round lobster cages, $1$ green seaweed patch, $1$ blue seaweed patch, $2$ tires, $1$ small hand net, $2$ types of towfish, $2$ pipes with anchors, $1$ star-shaped lobster cage
$4$
The vehicle position and heading are available at 1 Hz, and are directly used as the odometry input for the semantic SLAM method. The ground truth trajectory is further refined through a pose-graph SLAM method that incorporates ICP-based alignments of the wall and structures. The use of an outdoor tank on a sunny day means that the data has several difficult lighting effects, as seen in Figure 5. The objects in the tank are a fish cage, a ladder, a pipe, and a T-slot extrusion structure. There are also periodic wall markings from algae buildup and construction artifacts. The dataset consists of a single run of three loops around the circular tank, which is roughly 64 square meters in area with an average depth of 4 meters.
IV-B Indoor Pool
Data was also collected in an indoor tank with a greater variety of objects (Table I), but no available ground truth trajectory. The vehicle used for these experiments is a different modified BlueROV2, which is described in detail in [49]. Odometry is obtained by fusing a Teledyne Impulse Doppler velocity log, a VectorNav IMU, and a Bar30 pressure sensor using an Extended Kalman Filter. The camera is the same Sony Exmor IMX322 as with the outdoor dataset. The sonar is a Oculus m1200d multibeam operating at 2.1MHz, which provides a 2.5 mm range resolution, a 0.4 degree angular resolution, a 60 degree horizontal aperture, and a 12 degree vertical aperture. Despite the comparatively smaller vertical aperture, the camera data is still needed to constrain the object elevation angles.
@@ V Results
The proposed method is thoroughly tested on a variety of collected datasets, with map accuracy results shown in Table II, and trajectory accuracy along with an ablation study in Table III. Quantitative trajectory evaluations are available for the outdoor tank, but qualitative improvements of the trajectory can be seen in the indoor pool data as well in Figure 6. While the start and end point should be the same, the effects of odometric drift can be clearly seen as the end point of the odometry only trajectories are several meters from the start point. In contrast, our method is able to make loop closures through successfully re-identifying and associating objects as the vehicle returns to the start point, thus allowing the tightly-coupled optimization to correct the odometric drift. The outdoor tank dataset demonstrates the failure of feature-based methods in such conditions, as the standard implementation for such methods, ORBSLAM3 [50] is unable to complete more than $27\%$ of the trajectory due to the dynamic movement of caustics throughout the trajectory. Furthermore, our method is able to improve the trajectory accuracy from the odometry through identifying previously seen objects. The results also emphasize the success of the use of the semantic encoding of each object (e.g., Figure 7), as the ablation demonstrates that the trajectory is made worse without the assistance of the semantic latent codes due to incorrect data associations.
@@
Outdoor Tank UMAP (ours) 0.66 0.75

### v2

@@ I Introduction
Figure 2: The pipeline receives sensor inputs from a multibeam sonar, a monocular camera, an IMU, a DVL, and a pressure sensor. Given those sensor inputs, the system does the following: (1) Segment objects from optical images and extract features for each segmentation mask. The dimension of these extracted features is dependent on the corresponding object’s size in the image. These features are projected into a fixed dimension for the data association process. (2) Given the segmentation mask, its pixel centroid is used to find a corresponding range from sonar returns and estimate a 3D position of the object. (3) In parallel, sensor readings from IMU, DVL, and pressure sensor are used to estimate odometry. Outputs from (1)-(3) are used to build a factor graphwhich is optimized via iSAM2 to produce maximum a posteriori (MAP) estimates. Lastly, we obtain map and trajectory estimates as an output.
@@ II-B Opti-Acoustic Correspondences
Unfortunately, each sensor modality has a degeneracy stemming from the loss of the elevation angle for sonar, and the loss of the range measurement for monocular cameras. Some works seek to address the degeneracy directly for sonar-only mapping by learning the elevation angle [21], through the fusion of orthogonally mounted sonars [22], or through carefully handling the degeneracy in the back-end [23]. Stereo or multi-view for optical camera-only mapping has been explored as well, but suffers from issues with lighting conditions, reflections, caustics, and featureless areas such as plain walls [7]. However, combining the two sensor modalities to address the degeneracies of each sensor modality is likely the best way to fully utilize their complimentary geometries.
@@
Jang et al. [28] attempt to learn opti-acoustic feature correspondences through the use of a convolutional neural network, while Jang et al. [29] try to calculate opti-acoustic feature matching by applying a style-transfer method and through integration with SuperGlue for multi-session or multi-vehicle mapping. Lanthier et al. [7] combine a stereo camera with sonar and infrared sensors to perform mapping, but note that stereo cameras struggle with obtaining range measurements for underwater vision due to uniform and featureless regions, and difficult lighting effects such as reflections and caustics. Yang et al. [30] use sonar to provide an absolute scale to monocular odometry. Kim et al. [31] separately perform visual SLAM and sonar space carving in two different mapping modules, before combining the two for dense reconstruction. Rahman et al. [32] combine sonar and monocular camera data in a SLAM framework they refer to as SVIN2, but their mapping technique uses point features rather than objects.
The only work that performs fusion at the object level as required by our method, rather than at the feature level, is by Raaj et al. [33], who use a monocular camera and sonar in a particle filter, albeit only for a single object.
@@ III-B Opti-acoustic Fusion for 3D Projection
Refer to captionRefer to captionRefer to captionRefer to captionRaw ImageSegmented ObjectRaw SonarEnhanced for VisualizationTrue RangeErroneous Range
@@ III-C Semantic SLAM
For each existing landmark that is determined to meet a heuristically chosen similarity threshold, the Mahalanobis distance is calculated between the observed landmark and the existing landmark following the method described by Kaess et al. [47]. The distance follows a six degree-of-freedom chi-square distribution, and thus we use the test below where $l$ is the new landmark, $l_{j}$ is the existing landmark, and $\Sigma$ is the recovered covariance matrix.
@@
Figure 5: (Top): The outdoor tank used for data collection exhibits challenging lighting conditions, including reflections (left) and caustics (right). (Bottom): Selected raw (uncalibrated) frames from the vehicle’s monocular camera. These images demonstrate the caustics (ripple lighting effect due to surface waves), reflections, and sudden lighting changes make processing underwater imagery particularly difficult compared to terrestrial scenes. Feature-based methods fail to maintain tracking due to the dynamic textures; our method uses objects instead and proves to be robust to such lighting effects.
@@ IV Datasets
Figure 6: Comparison results of trajectory estimates, where the ’O’ and ’X’ markers denote the start and end points of each lap, respectively. The start and end points of the true trajectory are the same location, but drift causes the vehicle’s odometric estimate to deviate over time. (a) The two lap trajectory (Run 3) estimated by our patch averaging method (PA) can reduce drift errors from the odometry trajectory. (b) A similar comparison over four laps (Run 4), highlighting corrections made by our method against the odometry. Motion in the z-direction is negligible and thus is not shown.
@@
Refer to captionRefer to captionLobster cageSeaweedHand netTowfishTirePipe
@@
IV-A Indoor Pool
Data was collected in an indoor tank with a large variety of objects (Run 1-4 in Table I), but no available ground truth trajectory. The vehicle used for these experiments is a modified BlueROV2, which is described in detail in [49]. Odometry is obtained by fusing a Teledyne Impulse DVL, a VectorNav IMU, and a Bar30 pressure sensor using an Extended Kalman Filter. The camera is a calibrated Sony Exmor IMX322 low-light camera with an 80-degree horizontal field-of-view and a 64-degree vertical field-of-view. The sonar is a Oculus m1200d multibeam operating at 2.1MHz, which provides a 2.5 mm range resolution, a 0.4 degree angular resolution, a 60 degree horizontal aperture, and a 12 degree vertical aperture. Due to the vertical aperture, camera data is needed to constrain the objects’ elevation angles.
IV-B Outdoor Tank
Data was also collected in an outdoor tank on a sunny day with a modified BlueROV2 that has an acoustic array for the iLBL-USBL measurements attached to the top of the vehicle. The vehicle also has a Bar30 pressure sensor for depth measurements, the same camera as in the indoor datasets, and an Oculus m750d multibeam sonar operating at 750kHz frequency with a 5 Hz update rate. In this mode, the sonar has a range resolution of 4 mm, an angular resolution of 1 degree, a horizontal field-of-view of 130 degrees, and a vertical aperture of 20 degrees. The relatively large vertical aperture, which cannot be resolved in the received data, means that the use of the camera for estimating the elevation angle of objects is crucial.
@@ IV-B Outdoor Tank
Since magnetometer-based heading measurements are adversely affected by the magnetic interference of the tank walls and other nearby equipment, to get a more accurate ground truth trajectory, the vehicle heading is directly inferred from beamforming on the measurements from the vehicle’s acoustic array by solving for vehicle heading $\theta_{k}$:
@@
TABLE I: Summary of Datasets
@@
1 $1$ small hand net, $1$ large hand net, $2$ sets of stairs, $2$ recycling bins, $2$ types of towfish, $2$ trash bins, $2$ paddles $2$
2 $1$ model submarine, $1$ green seaweed patch, $1$ pipe, $1$ towfish, $1$ star-shaped lobster cage, $1$ tire, $1$ small hand net $2$
3 $2$ round lobster cages, $1$ green seaweed patch, $1$ blue seaweed patch, $2$ tires, $1$ small hand net, $2$ types of towfish, $2$ pipes with anchors, $1$ star-shaped lobster cage $2$
4 $2$ round lobster cages, $1$ green seaweed patch, $1$ blue seaweed patch, $2$ tires, $1$ small hand net, $2$ types of towfish, $2$ pipes with anchors, $1$ star-shaped lobster cage $4$
5 $1$ fish cage, $1$ ladder, $1$ pipe, $1$ T-slot extrusion structure $3$
The vehicle position and heading are available at $1$ Hz, and are directly used as the absolute (as opposed to relative as in the indoor datasets) odometry input for the semantic SLAM method. The ground truth trajectory is further refined through a pose-graph SLAM method that incorporates ICP-based alignments of the wall and structures. The use of an outdoor tank on a sunny day means that the data has several difficult lighting effects, as seen in Figure 5. The dataset consists of a single run (Run 5 in Table I) of three loops around the circular tank, which is roughly 64 square meters in area with an average depth of 4 meters. Note that while certain markings on the wall due to algae build-up were detected by the open-set system, those objects are not included the map accuracy results in Table II.
@@ V Results
The proposed method is thoroughly tested on a variety of collected datasets, with map accuracy results shown in Table II, and trajectory accuracy along with an ablation study in Table III. Quantitative trajectory evaluations are available for the outdoor tank, but qualitative improvements of the trajectory can be seen in the indoor pool data as well in Figure 6(b). While the start and end point should be the same, the effects of odometric drift can be clearly seen as the end points of the odometry only trajectories are several meters from the start point. In contrast, our method is able to make loop closures through successfully re-identifying and associating objects as the vehicle returns to the start point, thus allowing the tightly-coupled optimization to correct the odometric drift.
The outdoor tank dataset demonstrates the failure of feature-based methods in such conditions, as the standard implementation for such methods, ORB-SLAM3 [50] is unable to complete more than $27\%$ of the trajectory due to the dynamic movement of caustics throughout the trajectory. Furthermore, our method is able to improve the trajectory accuracy from the odometry through identifying previously seen objects. The results also emphasize the success of the use of the semantic encoding of each object (e.g., Figure 7), as the ablation demonstrates that the trajectory is made worse without the assistance of the semantic latent codes due to incorrect data associations.
@@
Outdoor Tank, Run 5 UMAP (ours) 0.66 0.75